OpenStudy (anonymous):

Let u = <-4, -3>. Find the unit vector in the direction of u, and write your answer in component form.

OpenStudy (bahrom7893):

I'm going to write the unit vector as ^u (u hat)

OpenStudy (bahrom7893):

First find L: L = sqrt((-4)^2 + (-3)^2) = sqrt(16+9) = sqrt(25) = 5

OpenStudy (anonymous):

is that a formula?

OpenStudy (bahrom7893):

Unit vector is simply: (1/L)*u = (1/5)*<-4,-3> = <-4/5; -3/5>

OpenStudy (bahrom7893):

yea the formula is basically: L = sqrt(x_comp^2 + y_comp^2 + x_comp^2 .. )

OpenStudy (anonymous):

thank you so much can you hep me on one more pls?

OpenStudy (anonymous):

Given that P = (5, 9) and Q = (13, 12), find the component form and magnitude of PQ (arrow on top)

OpenStudy (bahrom7893):

If I remember how to do those.

OpenStudy (bahrom7893):



OpenStudy (anonymous):

i have to use the formula right?

OpenStudy (bahrom7893):

well now we have to find the magnitude and the direction of the vector. Magnitude is pretty simple, it's just the length: